OpenStudy (anonymous):

Explain how to simplify the ratio x^2 - x - 20 over x^2 + 8x + 16

OpenStudy (diyadiya):

factorise x^2 - x - 20 and x^2 + 8x + 16

OpenStudy (diyadiya):

(x-5)(x+4)/(x+4)(x+4)

OpenStudy (diyadiya):

=(x-5)/(x+4)

OpenStudy (anonymous):

x^2-x-20=(x-5)(x-4) x^2 + 8x + 16=(x+4)^2 Divide both (x-5)/(x+4)

OpenStudy (anonymous):

thank you!